In the United States, the number of deaths from all causes was 70,663,474 for Whites, 9,692,906 for Blacks, and 1,125,108 for other racial groups for the combined years 2002-2006. The number of cancer deaths during this same time period was 15,533,648 for Whites, 1,946,813 for Blacks, and 249,591 for other racial groups. In which racial group is the proportional mortality ratio lowest? a. Whites b. Blacks c. Other
Answer #1

Answer :

Is " b " that is in Blacks.

Because proportional mortality ratio formula is,

PMR = Mortality due to a specific cause during a time period / Mortality due to all causes during in the same time period × 100.

So here according to that formula

PMR in whites = 15533648/70663474 ×100

= 21.98

PMR In blacks = 1946813/9692906 ×100

= 20.08

PMR in others = 249591/1125108 ×100

= 22.18

So the proportional mortality ratio lowest is in blacks that is 20.08.

Answer is b.